State the mathematical relationship between linear attenuation and mass attenuation co efficient
1
Expert's answer
2020-10-08T18:14:26-0400

According to Beer-Lambert law,

"I = I_0 e^{-\mu z} = I_0 e^{-\frac{\mu}{\rho}\lambda} = I_0 e^{-\alpha \lambda}"

where "I" is intensity, "\mu" - linear attenuation coefficient, "\alpha" - mass attenuation coefficient, "\lambda = \rho l" - the area density known also as mass thickness.

So, "\displaystyle \alpha = \frac{\mu}{\rho}".

Answer: "\displaystyle \alpha = \frac{\mu}{\rho}" .